# Zero-Knowledge Federated Learning: A New Trustworthy and Privacy-Preserving Distributed Learning Paradigm

### v2

@@ V-A Simulation Settings
We conduct numerical experiments based on the MNIST image dataset, with an FL process involving 20 clients, in which each client has 3000 sample data injected with Gaussian noise, at 4 different levels. Meanwhile, the server $S$ possesses a clean and small dataset of 400 sample data as its root dataset. During training, the used loss function is the symmetric cross entropy, which combines cross-entropy and inverse cross-entropy to address the over-fitting problem. As discussed in last section, the top $N$ local models are selected for aggregation after their performances’ ZKP is verified. Therefore, for the aggregation process, we set $N$ as 4, 8, 12, 16, and 20 respectively.
To investigate the computation complexity of ZKP generation and verification, and the size of proof, we set the model size (the length of the vectors $w^{t}_{k}$) to values from the set $\{8,16,32,64,128,256,512,1024,2048,4096\}$. The zk-SNARK algorithm of Groth16 [14] is used to generate and verify ZKP in Veri-CS-FL. We implement the Groth16 program for Veri-CS-FL using the library of Zokrates [15].
@@ V-B Results Analysis
We then investigate the performance of client selection in FL by training a model with $10^{3}$ parameters. In Fig. 2, we illustrate the accuracies of the global models aggregated in Veri-CS-FL under varying numbers of selected clients. The results demonstrate that as the number of selected clients decreases from 20 to 4, the global model’s performance consistently improves, peaking at over 90%. This improvement is likely due to the noise present in individual clients’ models—aggregating a larger number of noisy models can degrade overall performance. To benchmark the client selection performance of Veri-CS-FL, we also simulate an alternative FL scheme, Rand-CS-FL, where the server randomly selects $N$ clients from all the client to submit their local models, rather than selecting clients based on their models’ performance. In Fig. 2, we compare the performances of Veri-CS-FL and Rand-CS-FL with $N=4$. The results show that Veri-CS-FL outperforms Rand-CS-FL by effectively identifying clients with high-performing local models for aggregation, thereby improving the global model’s overall performance.
@@ VI Conclusion
In this article, we construct a systematic ZK-FL framework that classifies and examines the technical roles of ZKP across various FL modules. Through this framework, we provide a structured analysis of current ZK-FL approaches, identifying key areas where ZKP enhances both trust and efficiency in FL systems. Additionally, we propose the Verifiable Client Selection FL (Veri-CS-FL) algorithm, a novel approach that leverages ZKP to verify local model metrics based on their similarity to a benchmark model. By integrating ZKP, Veri-CS-FL ensures the reliable selection of high-quality models, improves model aggregation, and fosters trust among participants.

### v3

@@ V-A Simulation Settings
We conduct numerical experiments based on two image datasets: MNIST and FEMNIST. The MNIST dataset is used to evaluate Veri-CS-FL under independent and identically distributed (IID) settings, while the FEMNIST dataset, which is inherently partitioned by writer, is employed to evaluate the algorithm under non-IID settings, reflecting more realistic heterogeneous data distributions across clients. The FL process involves 20 clients, in which each client has 3000 sample data injected with Gaussian noise at 4 different levels. Meanwhile, the server $S$ possesses a clean and small dataset of 400 sample data as its root dataset. During training, the used loss function is the symmetric cross entropy, which combines cross-entropy and inverse cross-entropy to address the over-fitting problem. As discussed in last section, the top $N$ local models are selected for aggregation after their performances’ ZKP is verified. Therefore, for the aggregation process, we set $N$ as 4, 8, 12, 16, and 20 respectively.
To investigate the computation complexity of ZKP generation and verification, and the size of proof, we set the model size (the length of the vectors $w^{t}_{k}$) to values from the set $\{8,16,32,64,128,256,512,1024,2048,4096\}$. The zk-SNARK algorithm of Groth16 [4] is used to generate and verify ZKP in Veri-CS-FL. We implement the Groth16 program for Veri-CS-FL using the library of ZoKrates (https://zokrates.github.io/). All experiments were conducted on a server equipped with an Intel Xeon CPU and 64GB RAM. The reported proof generation and verification times include the core proving time of the Groth16 backend, where ZoKrates compiles the high-level code into R1CS (Rank-1 Constraint System) constraints for measurement.
@@ V-B Results Analysis
We then investigate the performance of client selection in FL by training a model with $10^{3}$ parameters. In Fig. 2, we illustrate the accuracies of the global models aggregated in Veri-CS-FL under varying numbers of selected clients. The results demonstrate that as the number of selected clients decreases from 20 to 4, the global model’s performance consistently improves, peaking at over 90%. This improvement is likely due to the noise present in individual clients’ models—aggregating a larger number of noisy models can degrade overall performance. To benchmark the client selection performance of Veri-CS-FL, we also simulate an alternative FL scheme, Rand-CS-FL, where the server randomly selects $N$ clients from all the client to submit their local models, rather than selecting clients based on their models’ performance. We selected Rand-CS-FL as the baseline instead of aggregating all models to ensure a fair comparison under identical communication constraints (selecting top $N$ clients), reflecting realistic edge scenarios where bandwidth is limited. In Fig. 2, we compare the performances of Veri-CS-FL and Rand-CS-FL with $N=4$. The results show that Veri-CS-FL outperforms Rand-CS-FL by effectively identifying clients with high-performing local models for aggregation, thereby improving the global model’s overall performance.
To validate the robustness of Veri-CS-FL on non-IID data scenarios, we conducted additional experiments using the FEMNIST dataset, which is explicitly designed for FL in non-IID settings. The number of selected clients is $N=4$. The root dataset includes 400 sample data randomly sampled from the whole dataset. The results, reported in Fig. 2, confirm that our client selection strategy remains effective in such scenarios, successfully filtering out low-quality updates and improving convergence speed compared to random selection. Specifically, in the non-IID FEMNIST case, Veri-CS-FL significantly accelerates early convergence by using the benchmark to filter out conflicting gradients caused by data heterogeneity. In contrast, for the IID MNIST case, Veri-CS-FL primarily excels in the later fine-tuning stages by filtering out noisy updates to reach a higher accuracy ceiling.
Regarding scalability, Veri-CS-FL scales efficiently. Server-side verification of Groth16 proofs is constant-time ($\approx$ milliseconds) and embarrassingly parallel, allowing the server to handle thousands of clients. Client-side proof generation is performed locally, ensuring horizontal scalability as the number of clients increases.
In terms of communication overhead, a Groth16 proof is approximately 128 bytes, and a SHA-256 hash is 32 bytes, totaling a negligible $\approx$ 160 bytes per client. In contrast, preventing the upload of even a single poorly performing model (which can be megabytes or gigabytes in size) yields bandwidth savings that far outweigh the cost of transmitting proofs. Thus, the overhead is $O(1)$ while the savings are O(Model Size)O($\text${Model Size}). This reduction in communication overhead enhances the effectiveness of Veri-CS-FL in bandwidth-constrained scenarios (e.g., wireless edge networks) or high-security contexts, where efficient data transmission and minimized communication costs are essential. Moreover, rapid advancements in hardware acceleration for ZKPs (e.g., on mobile GPUs/NPUs) are making client-side proof generation increasingly feasible, further enhancing the practicality of Veri-CS-FL on resource-constrained devices.
@@ VI Conclusion
In this article, we construct a systematic ZK-FL framework that classifies and examines the technical roles of ZKP across various FL modules. Through this framework, we provide a structured analysis of current ZK-FL approaches, identifying key areas where ZKP enhances both trust and efficiency in FL systems. Additionally, we propose the Verifiable Client Selection FL (Veri-CS-FL) algorithm, a novel approach that leverages ZKP to verify local model metrics based on their similarity to a benchmark model. By integrating ZKP, Veri-CS-FL ensures the reliable selection of high-quality models, improves model aggregation, and fosters trust among participants. While this work primarily addresses the upstream communication bottleneck, future research may explore integrating downstream compression techniques or selective model broadcasting to further reduce the total communication overhead in the FL lifecycle.